OpenStudy (anonymous):

solve the system by substitution -x-y-z=-8, -4x+4y+5z=7, 2x+2z=4

OpenStudy (raden):

from the 3rd equation : 2x+2z=4, divided by 2 we get : x+z=2 or z = 2 - x now subtitute z = 2 - x into the 1st equation : -x-y-z=-8 -x-y-(2-x)=-8 -x-y-2+x=-8 -y-2 = -8 -y = -8 + 2 -y = -6 y = 6 subtitute y = 6 into the 1st equation : -x-y-z=-8 -x-6-z=-8 -x-z=-8+6 -x-z=-2 -x = -2 + z x = z - 2 again subtitute y = 6 and x = z - 2 into the 2nd equation : -4x+4y+5z=7 -4(z - 2) + 4(6) + 5z = 7 -4z + 8 + 24 + 5z = 7 z + 32 = 7 z = 7 - 32 z = -25 because x = z - 2 (we got z = -25) so, x = -25 - 2 = -27 so, the solution (x,y,z) = (-27,6,-25)

OpenStudy (raden):

looks i make mistake :) wait a second

OpenStudy (raden):

given 3 systems : -x-y-z=-8 ... (1) -4x+4y+5z=7 ... (2) 2x+2z=4 ... (3) from the 3rd equation : 2x+2z=4, divided by 2 we get : x+z=2 or z = 2 - x ... (let it is the 4th equation) now subtitute z = 2 - x into the 1st equation : -x-y-z=-8 -x-y-(2-x)=-8 -x-y-2+x=-8 cancel the x's, it can be -y-2 = -8 or -y = -8 + 2 -y = -6 y = 6 then subtitute y = 6 and z = 2 - x into the 2nd equation : -4x+4y+5z=7 -4x + 4(6) + 5(2 - x) = 7 -4x + 24 + 10 - 5x = 7 -9x + 34 = 7 -9x = 7 - 34 -9x = -27 x = -27/-9 x = 3 finally, subtitute x = 3 into the 4th equation : z = 2 - x z = 2 - 3 z = -1 so, the solution (x,y,z) = (3, 6, -1)

OpenStudy (anonymous):

thanks you so much ::))))))

OpenStudy (raden):

you're welcome dear :*